OpenStudy (anonymous):

Which literal equations are equivalent to k=f/x Choose all answers that are correct. a. f=kx b. x=kf c. f=k/x d. x=f/k

OpenStudy (anonymous):

\[k= \frac{ F }{ x }\] Multiplying both sides by x \[k*x = \frac{ F*x }{ x }\] So what will be the answer?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Correct :)

OpenStudy (anonymous):

and

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

Choose all answers that are correct.

OpenStudy (anonymous):

ok there is another option which will be selected. Sorry I missed that one. Do you know which one?

OpenStudy (anonymous):

c but im not sure thats correct

OpenStudy (anonymous):

is it

OpenStudy (anonymous):

Its not c Last equation we did was \[k*x = F\] Now we can divide both sides with k to find the value of x so the equation will be \[\frac{ k*x }{ k } = \frac{ F }{ k }\] \[x = \frac{ F }{ k }\]

OpenStudy (anonymous):

so d

OpenStudy (anonymous):

yes